Two particles each of rest mass m collide head-on and stick together. Before collision, the
speed of each mass was 0.6 times the speed of light in free space. The mass of the final
entity is
1
Expert's answer
2020-09-29T09:42:23-0400


"E_intial=E_final"


"Final velocity=0"


"\frac {MC^2}{\sqrt1-(0.6)^2}+\frac {MC^2}{\sqrt1-(0.6)^2}= \frac {MC^2}{\sqrt1-(0)}"



"\frac{2MC^2}{0.8}"



"M=\frac{5m}{2}"